OpenStudy (anonymous):

Sequence/recurrence relation. Find the value of t(5) if t(n+1) = 3t(n) + 2. Note the numerals in brackets are supposed to be little underneath numbers.

OpenStudy (anonymous):

\[t _{(n+1)}=3t _{(n)}+2\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

oh wait I can write it now with equation

OpenStudy (anonymous):

but how you you do it?

OpenStudy (anonymous):

do the problem or write out the equation?

OpenStudy (anonymous):

Do the problem. I don't get it really.

OpenStudy (anonymous):

i was first seeing if what i typed is how the problem looks like, is it correct? ^^^

OpenStudy (anonymous):

\[t _{5}\]

OpenStudy (anonymous):

yes that is right. That is the problem. It is a sequence equation.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what is \[t _{5}\]?

OpenStudy (anonymous):

Shouldn't \[t_{5}\] be 14? But he answer says no.

OpenStudy (anonymous):

that is the term. if \[t _{0}\] is the first term, \[t _{5}\] should be the sixth term

OpenStudy (anonymous):

Oh sorry \[t _{1}=1\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you're suppose to get an answer without variable t?

OpenStudy (anonymous):

\[t _{1} = 1\] Yeah I don't really get it. I think t is just really like the bit under the t.

OpenStudy (anonymous):

can you upload an image of the question?

OpenStudy (anonymous):

Not really capable. But that is all the question really says, excpet there is a whole tutorial section right before it, trouble is I don't understand it. There is an example kind of like this: if the rule is \[t _{1}= 1 and t _{k+1} = 2t _{k}\] then \[t _{1}=1, t _{2} = 2t _{1}= 2, t _{3}= 2t _{2}=\]4

OpenStudy (anonymous):

so \[t _{5} should be 4t _{4}\]

OpenStudy (anonymous):

why? I don't know. The answer is supposed to be 161

OpenStudy (anonymous):

so the first thing to do is find what \[t _{_{(n+1)}} is....t _{(5+1)}=?\]

OpenStudy (anonymous):

i was following the logic from your tutorial, it's set up the same way

OpenStudy (anonymous):

\[3t _{4} + 2\]

OpenStudy (anonymous):

no we're both wrong it's \[2t _{3}\]

OpenStudy (anonymous):

+2

OpenStudy (anonymous):

if you want to see step by step,\[t _{1}=1\] \[t _{2}=2t _{1}=2 t _{1}=1\] so just plug 1 in which would make it 2*1

OpenStudy (anonymous):

\[t _{3}=2t _{2}=4\] \[t _{4}=2t _{3}=8\] \[t _{5}=2t _{4}=16\]

OpenStudy (anonymous):

note that i didn't +2 in any one them...i just realized

OpenStudy (anonymous):

but just follow the same logic

OpenStudy (anonymous):

I'm getting there lol thankyou for the help!

OpenStudy (anonymous):

np

OpenStudy (mathmate):

t(1)=1 t(2)=3t(1)+2=5 t(3)=3t(2)+2=17 t(4)=3t(3)+2=53 t(5)=3t(4)+2=161

OpenStudy (anonymous):

thanks mathmate. I have now understanding of this.

OpenStudy (mathmate):

Good! Glad to be of help!